OpenStudy (anonymous):

I NEED HELP!! PLZ! WILL GIVE MEDAL AND FAN!!! solve the equation. check the solution 4=√p-2 A.√6 B.36 C.3 D.6 solve the quation identify any extraneous solutions. 8√9j+10=1 A.9/64 is a solution of the original equation. -9/64 is the ES B.9/64 is the solution of the original equation C.-9/64 is the solution of the original equation. 9/64 is an ES D.there is no solution

OpenStudy (anonymous):

\[4=\sqrt{p}-2\] \[4+2=\sqrt{p}\] \[6=\sqrt{p}\] \[6^2=(\sqrt{p})^2\] \[\boxed{\boxed{p=36}}\] ____________________________________________________________________ \[8*\sqrt{9j}+10=1\] \[8*\sqrt{9j}=1-10\] \[8*\sqrt{9j}=-9\] \[8*3\sqrt{j}=-9\] \[8\sqrt{j}=-3\] \[\sqrt{j}=-\frac{3}{8}\] \[(\sqrt{j})^2=\left(-\frac{3}{8}\right)^2\] \[\boxed{\boxed{j=\frac{9}{64}}}\]

OpenStudy (anonymous):

so for the second one would it be A orB?

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

B

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (anonymous):

what is the domain of the function? y=√4x-4 A.x>-1 B. x> or = to 1 C.x> or equal to -1 D.x> or equal to

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

the function 4x-4 never can be negative then: \[y=\sqrt{4x-4}\] \[4x-4\geq0\] \[\boxed{\boxed{x\geq1}}\]

OpenStudy (anonymous):

ok thanks. what about this one what is the cos for for the triangle shown A.4/3 B.4/5 C.5/4 D.3/5

OpenStudy (anonymous):

@D3xt3R

OpenStudy (anonymous):

please help! I will give u plenty of medals lol

OpenStudy (anonymous):

I have a test now, when I'll back i answer